maddie123:

Which equation is true for the value a = 7? A. a + 2(a – 3) = 15 B. 4a – 4(14 – a) = 24 C. 2a + 47 = 105 D. 3a + 2(a – 3) = 34 E. 4a - 2a = 18

bizaardvark:

If a = 7 then you would want to plug in 7 for a in each equation to see which one comes out to the right answer. First we can try A: a + 2(a – 3) = 15 (7) + 2 ( (7) - 3) =15 7 + 2 (4) = 15 7 + 8 = 15 15 = 15 Since 15 does equal 15 that tells us the answer is correct

maddie123:

it dose not match A with 7

bizaardvark:

what do you mean?

bizaardvark:

did you see the equation i did above??

bizaardvark:

If you plug 7 into any of the other equations, the answer would equal itself, so it would prove that those other options are incorrect

maddie123:

so its A

bizaardvark:

yup

bizaardvark:

do you understand it though?

maddie123:

ok thx

bizaardvark:

alr lol np :)